40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 10 15:41:37 2019, aaron, Update, IOO, WFS measurements Screenshot_from_2019-09-10_18-51-28.pngmctrend_190910.png
    Reply  Wed Sep 11 10:26:56 2019, aaron, Update, IOO, WFS measurements 
       Reply  Wed Sep 11 14:37:43 2019, aaron, Update, IOO, WFS measurements 20190911_WFS.jpg20190911_WFS_2.jpg
          Reply  Thu Sep 12 12:42:31 2019, aaron, Update, IOO, WFS measurements WFS2_TFs.pdfD755499D-9FDF-4E2B-BFC1-016B459DD35D.jpeg
             Reply  Fri Sep 13 10:36:03 2019, aaron, Update, IOO, WFS measurements TF-.pngWFS2_TFs.pdf
                Reply  Mon Sep 16 12:38:59 2019, aaron, Update, IOO, WFS measurements 
                Reply  Thu Oct 3 11:38:35 2019, aaron, Update, IOO, WFS measurements 
          Reply  Mon Sep 30 13:20:55 2019, aaron, Update, IOO, shot noise measurement 
       Reply  Fri Sep 13 10:53:40 2019, aaron, Update, IOO, WFS loop measurements 
          Reply  Mon Sep 16 05:08:04 2019, rana, Update, IOO, WFS loop measurements LSCsensingMatrix.ipynbASCsensingMatrix.ipynb
             Reply  Mon Sep 16 11:55:58 2019, rika, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 09:41:48 2019, gautam, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 10:47:44 2019, rika, Update, IOO, WFS loop measurements MC2.pdf
                   Reply  Wed Sep 18 14:45:52 2019, rika, Update, IOO, WFS loop measurements Screenshot_from_2019-09-18_18-15-34.png
             Reply  Tue Sep 17 10:34:48 2019, aaron, Update, IOO, WFS loop measurements 
                Reply  Tue Oct 8 20:39:42 2019, aaron, Update, IOO, WFS loop measurements 
                   Reply  Wed Oct 9 09:37:28 2019, aaron, Update, IOO, WFS loop measurements 
Message ID: 14872     Entry time: Wed Sep 11 14:37:43 2019     In reply to: 14871     Reply to this: 14874   14914
Author: aaron 
Type: Update 
Category: IOO 
Subject: WFS measurements 

[aaron, rika]

We identified the Jenne laser and found a long optical fiber that might be able to transport our beam to the AP table.

Now we're searching for documentation on using this laser. Kevin and John measured a TF last year. Koji advised that we needn't worry too much, the current limit is already set correctly and we need only power on the laser.

We moved the breadboard (including a couple PDs, collimating lenses, laser, steering mirrors, etc) over to the AP table, and set it on top of the panel next to the WFS. We mounted the laser on the AP table, and added one lens with f~68 mm after the laser to fit the beam on a single quadrant; the beam was about 1mm diameter (measured by eye) when it entered the QPD. We turned the laser driver on at ~19.4 mA, and directed it to WFS2 via the last two steering mirrors before WFS2.

We monitored the QPD segments' DC level with ndscope on a laptop, and were able to send the beam to each of the four quadrants in turn. We set up the Agilent network analyzer to drive the laser's amplitude modulation and sent the RF signal from the LEMO output on the QPD head directly to the network analyzer. We will take the measurements tomorrow morning.

Attachment 1: 20190911_WFS.jpg  201 kB  Uploaded Wed Sep 11 20:35:05 2019  | Hide | Hide all
20190911_WFS.jpg
Attachment 2: 20190911_WFS_2.jpg  246 kB  Uploaded Wed Sep 11 20:35:33 2019  | Hide | Hide all
20190911_WFS_2.jpg
ELOG V3.1.3-